OpenStudy (anonymous):

z^(-3)/z^(-7) how do I get rid of the negitive exponent

OpenStudy (anonymous):

Get the inverse of it. \[z^{(-3)}=\frac{1}{z^{3}} Likewise \frac{1}{z^{(-7)}}=z^{7}\]

OpenStudy (anonymous):

Remember that \(\Large\frac{x^{a}}{x^{b}} = x^{a - b}\) It doesn't matter if \(a\) and/or \(b\) are positive or negative. The standard rules of addition and subtraction apply.